OpenStudy (anonymous):

Question 4 If we toss 1 coin, what is the probability of getting heads? Answer 1/6 1/8 1/4 1/2 1 points Question 5 If we toss one die, what is the probability of getting a 4? Answer 1/6 1/4 1/2 1/3

OpenStudy (anonymous):

@ganeshie8

OpenStudy (mayankdevnani):



OpenStudy (mayankdevnani):

if we toss 1 coin,total outcomes =2 (head or tail )

OpenStudy (mayankdevnani):

getting head for tossing a coin 1 time = 1

OpenStudy (mayankdevnani):

so,\[\large \bf probability=\frac{favourable~outcomes}{total~outcomes}=\frac{1}{2}\]

OpenStudy (mayankdevnani):

understood ?? @MeggyMeg

OpenStudy (mayankdevnani):

for second question :- if we thrown a die,total outcomes =6 (1,2,3,4,5,6) favourable outcomes(getting 4)=1

OpenStudy (mayankdevnani):

so, \[\large \bf probability =\frac{favourabe~outcomes}{total~outcomes}=\frac{1}{6}\]

OpenStudy (mayankdevnani):

hope you understand ! @MeggyMeg

OpenStudy (mayankdevnani):

and also \[\large \bf Welcome~TO~OpenSTUDY\]

OpenStudy (anonymous):

ty!

OpenStudy (mayankdevnani):

welcome :)

OpenStudy (mayankdevnani):

you can also give me MEDAL by clicking `best response` flashing right side to my comment

OpenStudy (anonymous):

i just did :]

OpenStudy (mayankdevnani):

good !